# Limit theorems for a strongly irreducible product of independent random matrices under optimal moment assumptions

### v2

@@ Abstract.
If we moreover assume that $\nu$ is supported on the group of invertible matrices and that the push-forward distribution $N_{*}\nu$ is $\mathrm{L}^{p}$ for $N:g\mapsto\log\|g\|\|g^{-1}\|$ and for some $p>0$, then we show that $-\log\mathrm{d}(l^{\infty},H)$ is uniformly $\mathrm{L}^{p}$ for all proper subspace $H\subset\mathbb{R}^{d}$. For $p=1$, we moreover show that the rescaled logarithm of each coefficient of $\overline{\gamma}_{n}$ almost surely converges to the top Lyapunov exponent. This is an extension of results by Benoist and Quint which were themselves improvements of historic results by Guivarc’h, Lepage and Raugi, first proven under exponential moment assumptions.
@@ 1. Introduction
Let $d\geq 2$. To a random independent and identically distributed sequence $\left(\gamma_{n}\right)_{n\geq 0}\in\mathrm{GL}(\mathbb{R}^{d})$, we associate the random walk $\left(\overline{\gamma}_{n}\right)_{n\geq 0}$ defined as $\overline{\gamma}_{n}:=\gamma_{0}\cdots\gamma_{n-1}$ for all $n\geq 0$. The random matrix $\overline{\gamma}_{n}$ is then a product of $n$ independent random $d\times d$ matrices. We are interested in quantitative probabilistic estimates for $\overline{\gamma}_{n}$ that hold for all time $n$ as well as qualitative asymptotic results that hold almost surely for the sequence $\left(\overline{\gamma}_{n}\right)_{n\geq 0}$. We write ${}^{t}e_{i}\overline{\gamma}_{n}e_{j}$ for the entry of $\overline{\gamma}_{n}$ at the coordinates $(i,j)$. We write $\rho_{1}(\overline{\gamma}_{n})\geq\rho_{2}(\overline{\gamma}_{n})\geq\dots\geq\rho_{d}(\overline{\gamma}_{n})\geq 0$ for the spectral values of $\overline{\gamma}_{n}$ i.e., the moduli of the eigenvalues. We say that a matrix $h$ is proximal when $\rho_{1}(h)>\rho_{2}(h)$, which implies that the eigenspace of $h$ associated to its top eigenvalue is a line and that $\rho_{1}(h)^{-n}h^{n}$ converges to a rank one projection or equivalently that the projective class $[h^{n}]$ converges to the projective class of a rank one projection.
@@ 1.1. Preliminary definitions
[…]
@@
The condition (1) is trivially satisfied when $\nu$ is supported on the set $\mathrm{GL}(E)$ of invertible matrices. We show in Lemma 3.2 that we have a $0-1$ law for $\nu^{*n}\{0\}$ in the sense that we either have $\nu^{*n}\{0\}=0$ for all $n$ or there exists $C,\beta>0$ such that $\nu^{*n}\{0\}\geq 1-Ce^{-\beta n}$ for all $n\geq 0$. So when the second condition for proximality is not satisfied, the sequence $(\overline{\gamma}_{n})_{n\geq 0}$ is almost surely stationary to 00 for $(\gamma_{n})\sim\nu^{\otimes\mathbb{N}}$ and the random first time $n_{0}$ such that $\overline{\gamma}_{n_{0}}=0$ has a finite exponential moment.
@@
Note that condition (1) is not an algebraic condition. Indeed, if we consider $\nu$ to be a Gaussian distribution over $\mathrm{End}(E)$, then we have $\Gamma_{\nu}=\mathrm{End}(E)$ and $\nu(\mathrm{GL}(E))=1$ so $\nu$ satisfies conditions (2), (1), (3) and (4), hence it is strongly irreducible and proximal. On the other hand, if $\nu$ is the barycentre of a Gaussian distribution with the Dirac measure at 00, then we still have $\Gamma_{\nu}=\mathrm{End}(E)$, so conditions (2), (3) and (4) are satisfied, but $\nu^{*n}\{0\}=1-2^{-n}$ for all $n$ so (1) is not satisfied.
@@ 1.2. Alignment and pivotal extraction
Let $\mathbb{K}$ be a local field and let $d\geq 2$. Let $\nu$ be a strongly irreducible and proximal probability distribution over $\Gamma=\mathrm{Mat}_{d\times d}(\mathbb{K})=\mathrm{End}(\mathbb{K}^{d})$ or $\Gamma=\mathrm{GL}(\mathbb{K}^{d})$. There exist constants $20<\varepsilon\leq 1/$ 2 and $m\in\mathbb{N}$, a compact $K\subset\Gamma$ and a probability distribution $\mu$ on $\Gamma^{\mathbb{N}}\times\mathbb{N}\setminus\{0\}^{\mathbb{N}}$ such that given $\left(\left(\gamma_{n}\right)_{n\geq 0},\,\left(p_{n}\right)_{n\geq 0}\right)\sim\mu$, the following assertions hold.
@@ 1.3. Contraction results without moment assumptions
We denote by $\mathrm{P}(E)$ the projective space associated to $E$ i.e., the set of vector lines in $E$. Write $[\cdot]:E\setminus\{0\}\to\mathrm{P}(E)$ for the projection map. We endow $\mathrm{P}(E)$ with the metric:
@@
When $v$ is such that $\sup_{n}\mathbb{P}(\overline{\gamma}_{n}v=0)=0$ (which is true for $v$ in a dense and uncountable subset of $\mathrm{P}(E)$ by Lemma 3.4) (8) implies that $[\overline{\gamma}_{n}v]\to l^{\infty}$ almost surely.
@@
Let $0<\alpha<\lambda_{1-2}(\nu)$ and let $C,\beta>0$ be constants that satisfy (8) in Theorem 1.8. Let $[x],({\gamma}_{n})_{n}\sim\xi\otimes\nu^{\otimes\mathbb{N}}$. Since $\mathbb{P}(x\in\underline{\ker}(\nu))=0$, we have $\overline{\gamma}_{n}x\neq 0$ almost surely and for all $n$. Let $l^{\infty}$ be as in Theorem 1.8. Then $\mathbb{E}(f(l^{\infty}))=\int f\mathrm{d}\xi_{\nu}^{\infty}$ for all $f$. Moreover, for all $n$, we have $\mathbb{P}\left(\mathrm{d}([\overline{\gamma}_{n}x],l^{\infty})\leq e^{-\alpha n}\right)\leq Ce^{-\beta n}$. Moreover the distance map takes values in $[0,1]$ so $\mathbb{E}\left(\mathrm{d}([\overline{\gamma}_{n}x],l^{\infty})\right)\leq Ce^{-\beta n}+e^{-\alpha n}\leq C^{\prime}e^{-\beta^{\prime}n}$ for $C^{\prime}=C+1$ and $\beta^{\prime}=\min\{\beta,\alpha\}$. Then, by convexity of the absolute value, we have $\left|\mathbb{E}\left(f[\overline{\gamma}_{n}x]\right)-\mathbb{E}\left(f(l^{\infty})\right)\right|\leq\mathbb{E}\left(|f[\overline{\gamma}_{n}x]-f(l^{\infty})|\right)\leq\mathbb{E}\left(\lambda(f)\mathrm{d}([\overline{\gamma}_{n}x],l^{\infty})\right)\leq\lambda(f)C^{\prime}e^{-\beta^{\prime}n}$. ∎
Note that saying that $\xi$ is supported on $\mathrm{P}(E)\setminus\underline{\ker}(\nu)$ is not very restrictive because any measure that gives measure 00 to all hyperplanes would satisfy that condition. However, $\xi_{\nu}^{\infty}$ itself may give positive measure to some hyperplanes. For example if $\nu$ is the barycentre of the Haar measure on the group of isometries and a Dirac mass $\delta_{\pi}$ at a projection endomorphism $\pi$, then $\xi^{\infty}_{\nu}$ is the average of the isometry-invariant measure and of the Dirac mass on the image of $\pi$. In particular $\xi^{\infty}_{\nu}$ gives positive measure to any hyperplane that contains the image of $\pi$.
@@
[…]
@@ 1.5. Regularity results with optimal moment assumptions
The term $Ce^{-\beta n}$ in (11) naturally appears in the proof. This term is annoying because it prevents us from giving a bound on the moments of $-\log\frac{|f\overline{\gamma}_{n}v|}{\|f\|\|\overline{\gamma}_{n}\|\|v\|}$. However on can not get rid of it, to give a clear example, let $\nu$ be a lazy measure that gives measure $1/2$ to the identity matrix, then given $fv=0$, we have $\mathbb{P}(f\overline{\gamma}_{n}v=0)\geq 1-2^{-n}$ for all $n$.
@@
Note that $\sum_{k=1}^{+\infty}C\exp(-\beta k)<+\infty$ so $\sum_{k=1}^{+\infty}C\exp(-\beta k)N_{*}{\nu}\left(t/k,+\infty\right)$ goes to 00 as $t\to+\infty$ by the monotonous convergence Theorem. Moreover, since everything is non-negative, we have:
[…]
@@
[…]
@@
[…]
@@
Let $l^{\infty}$ be as in Theorem 1.8. We have $l^{\infty}\sim\xi_{\nu}^{\infty}$ so $\xi_{\nu}^{\infty}(\mathcal{N}_{r}(V))=\mathbb{P}(l^{\infty}\in\mathcal{N}_{r}(V))$. Moreover, by Theorem 1.8, we have $[\overline{\gamma}_{n}u]\to l^{\infty}$ almost surely and $\mathcal{N}_{r}(V)$ is open, so assuming that $l^{\infty}\in\mathcal{N}_{r}(V)$ implies that there exists $n_{0}$ such that $[\overline{\gamma}_{n}u]\in\mathcal{N}_{r}(V)$ for all $n\geq n_{0}$. Therefore, we have $\xi_{\nu}^{\infty}(\mathcal{N}_{r}(V))\leq\lim_{n}\mathbb{P}([\overline{\gamma}_{n}u]\in\mathcal{N}_{r}(V))$.
@@ 2.1. Alignment and contraction coefficient
We remind that we defined the notation $\mathbb{A}^{\varepsilon}$ in Definition 1.3, given $g$ and $h$ two matrices such that the product $gh$ is well defined (i.e., $g$ has as many columns as $h$ has rows) and $10<\varepsilon<$ 1, we write $g\mathbb{A}^{\varepsilon}h$ as a short notation for $\|gh\|\geq\varepsilon\|g\|\|h\|$. Note that $g$ may have a single line and $h$ may have a single column, in which case $gh$ is a $1\times 1$ square matrix i.e., a scalar. We remind the reader that given $x,y$ two vectors, we have the characterization $\|x\wedge y\|=\min_{a\in\mathbb{K}}\|x-ya\|\|y\|$. Therefore, given $h\in\mathrm{Hom}(E,F)$, we have $\|h\wedge h\|=\max_{x,y\in E\setminus\{0\}}\min_{a\in\mathbb{K}}\frac{\|h(x-ya)\|\|h(y)\|}{\|x\|\|y\|}$.
@@
Note also that for $h$ an endomorphism of rank one, and for all $0<\varepsilon\leq 1$, the cone $U^{\varepsilon}(h)$ is the image of $h$ so it has diameter 00 in the projective space.
@@
The second fact is that (${}^{t}(gh)=^{t}h^{t}g$. It implies that for all non trivial $g,h$, we have $\frac{\|gh\|}{\|g\|\|h\|}=\frac{\|{{}^{t}h}{{}^{t}g}\|}{\|{{}^{t}h}\|{{}^{t}g}\|}$. The third fact is that ${}^{t}h\wedge^{t}h=^{t}(h\wedge h)$. It implies that $\sigma(^{t}h)=\sigma(h)$. ∎

### v3

@@ Abstract.
If we moreover assume that $\nu$ is supported on the group of invertible matrices and that the push-forward distribution $N_{*}\nu$ is $\mathrm{L}^{p}$ for $N:g\mapsto\log\|g\|\|g^{-1}\|$ and for some $p>0$, then we show that $\log\mathrm{d}(l^{\infty},H)$ is uniformly $\mathrm{L}^{p}$ for all proper subspace $H\subset\mathbb{R}^{d}$. For $p=1$, we moreover show that the rescaled logarithm of each coefficient of $\overline{\gamma}_{n}$ almost surely converges to the top Lyapunov exponent. This is an extension of results by Benoist and Quint which were themselves improvements of historic results by Guivarc’h, Lepage and Raugi, first proven under exponential moment assumptions.
@@ 1. Introduction
Let $d\geq 2$. To a random independent and identically distributed sequence $\left(\gamma_{n}\right)_{n\geq 0}\in\mathrm{GL}(\mathbb{R}^{d})$, we associate the random walk $\left(\overline{\gamma}_{n}\right)_{n\geq 0}$ defined as $\overline{\gamma}_{n}:=\gamma_{0}\cdots\gamma_{n-1}$ for all $n\geq 0$. The random matrix $\overline{\gamma}_{n}$ is then a product of $n$ independent random $d\times d$ matrices. We are interested in quantitative probabilistic estimates for $\overline{\gamma}_{n}$ that hold for all time $n$ as well as qualitative asymptotic results that hold almost surely for the sequence $\left(\overline{\gamma}_{n}\right)_{n\geq 0}$. We write $^{t}e_{i}\overline{\gamma}_{n}e_{j}$ for the entry of $\overline{\gamma}_{n}$ at the coordinates $(i,j)$. We write $\rho_{1}(\overline{\gamma}_{n})\geq\rho_{2}(\overline{\gamma}_{n})\geq\dots\geq\rho_{d}(\overline{\gamma}_{n})\geq 0$ for the spectral values of $\overline{\gamma}_{n}$ i.e., the moduli of the eigenvalues. We say that a matrix $h$ is proximal when $\rho_{1}(h)>\rho_{2}(h)$, which implies that the eigenspace of $h$ associated to its top eigenvalue is a line and that $\rho_{1}(h)^{-n}h^{n}$ converges to a rank one projection or equivalently that the projective class $h^{n}]$ converges to the projective class of a rank one projection.
@@ 1.1. Preliminary definitions
Given $1\leq k\leq d$, we write $\bigwedge^{k}E$ for the $k$-th exterior product of $E$, which is simply the dual space of the space of $k$-linear alternate forms. For all $j,k\geq 1$ such that $j+k\leq d$, we have a bi-linear map $\wedge:\bigwedge^{j}E\times\bigwedge^{k}E\to\bigwedge^{j+k}E$. There is also a canonical norm $\|\cdot\|$ on $\bigwedge^{k}E$ for all $k$, which is characterized by the fact that for all $j,k$, for all $x\in\bigwedge^{j}E$ and for all $y\in\bigwedge^{k}E$, we have $\|x\wedge y\|\leq\|x\|\|y\|$ and this inequality is sharp in the sense that there exists $y^{\prime}\in\bigwedge^{k}E$ such that $x\wedge y=x\wedge y^{\prime}$ and such that $\|x\wedge y^{\prime}\|=\|x\|\|y^{\prime}\|$.
@@
The condition (1) is trivially satisfied when $\nu$ is supported on the set $\mathrm{GL}(E)$ of invertible matrices. We show in Lemma 3.2 that we have a 010-1 law for $\nu^{*n}\{0\}$ in the sense that we either have $\nu^{*n}\{0\}=0$ for all $n$ or there exists $C,\beta>0$ such that $\nu^{*n}\{0\}\geq 1-Ce^{-\beta n}$ for all $n\geq 0$. So when the second condition for proximality is not satisfied, the sequence $(\overline{\gamma}_{n})_{n\geq 0}$ is almost surely stationary to 0 for $(\gamma_{n})\sim\nu^{\otimes\mathbb{N}}$ and the random first time $n_{0}$ such that $\overline{\gamma}_{n_{0}}=0$ has a finite exponential moment.
@@
Note that condition (1) is not an algebraic condition. Indeed, if we consider $\nu$ to be a Gaussian distribution over $\mathrm{End}(E)$, then we have $\Gamma_{\nu}=\mathrm{End}(E)$ and $\nu(\mathrm{GL}(E))=1$ so $\nu$ satisfies conditions (2), (1), (3) and (4), hence it is strongly irreducible and proximal. On the other hand, if $\nu$ is the barycentre of a Gaussian distribution with the Dirac measure at 0, then we still have $\Gamma_{\nu}=\mathrm{End}(E)$, so conditions (2), (3) and (4) are satisfied, but $\nu^{*n}\{0\}=1-2^{-n}$ for all $n$ so (1) is not satisfied.
@@ 1.2. Alignment and pivotal extraction
Let $\mathbb{K}$ be a local field and let $d\geq 2$. Let $\nu$ be a strongly irreducible and proximal probability distribution over $\Gamma=\mathrm{Mat}_{d\times d}(\mathbb{K})=\mathrm{End}(\mathbb{K}^{d})$ or $\Gamma=\mathrm{GL}(\mathbb{K}^{d})$. There exist constants $0<\varepsilon\leq 1/2$ and $m\in\mathbb{N}$, a compact $K\subset\Gamma$ and a probability distribution $\mu$ on $\Gamma^{\mathbb{N}}\times\mathbb{N}\setminus\{0\}^{\mathbb{N}}$ such that given $\left(\left(\gamma_{n}\right)_{n\geq 0},\,\left(p_{n}\right)_{n\geq 0}\right)\sim\mu$, the following assertions hold.
@@ 1.3. Contraction results without moment assumptions
We denote by $\mathrm{P}(E)$ the projective space associated to $E$ i.e., the set of vector lines in $E$. Write $\cdot]:E\setminus\{0\}\to\mathrm{P}(E)$ for the projection map. We endow $\mathrm{P}(E)$ with the metric:
@@
When $v$ is such that $\sup_{n}\mathbb{P}(\overline{\gamma}_{n}v=0)=0$ (which is true for $v$ in a dense and uncountable subset of $\mathrm{P}(E)$ by Lemma 3.4) (8) implies that $\overline{\gamma}_{n}v]\to l^{\infty}$ almost surely.
@@
Let $0<\alpha<\lambda_{1-2}(\nu)$ and let $C,\beta>0$ be constants that satisfy (8) in Theorem 1.8. Let $x],({\gamma}_{n})_{n}\sim\xi\otimes\nu^{\otimes\mathbb{N}}$. Since $\mathbb{P}(x\in\underline{\ker}(\nu))=0$, we have $\overline{\gamma}_{n}x\neq 0$ almost surely and for all $n$. Let $l^{\infty}$ be as in Theorem 1.8. Then $\mathbb{E}(f(l^{\infty}))=\int f\mathrm{d}\xi_{\nu}^{\infty}$ for all $f$. Moreover, for all $n$, we have $\mathbb{P}\left(\mathrm{d}([\overline{\gamma}_{n}x],l^{\infty})\leq e^{-\alpha n}\right)\leq Ce^{-\beta n}$. Moreover the distance map takes values in $[0,1]$ so $\mathbb{E}\left(\mathrm{d}([\overline{\gamma}_{n}x],l^{\infty})\right)\leq Ce^{-\beta n}+e^{-\alpha n}\leq C^{\prime}e^{-\beta^{\prime}n}$ for $C^{\prime}=C+1$ and $\beta^{\prime}=\min\{\beta,\alpha\}$. Then, by convexity of the absolute value, we have $\left|\mathbb{E}\left(f[\overline{\gamma}_{n}x]\right)-\mathbb{E}\left(f(l^{\infty})\right)\right|\leq\mathbb{E}\left(|f[\overline{\gamma}_{n}x]-f(l^{\infty})|\right)\leq\mathbb{E}\left(\lambda(f)\mathrm{d}([\overline{\gamma}_{n}x],l^{\infty})\right)\leq\lambda(f)C^{\prime}e^{-\beta^{\prime}n}$. ∎
Note that saying that $\xi$ is supported on $\mathrm{P}(E)\setminus\underline{\ker}(\nu)$ is not very restrictive because any measure that gives measure 0 to all hyperplanes would satisfy that condition. However, $\xi_{\nu}^{\infty}$ itself may give positive measure to some hyperplanes. For example if $\nu$ is the barycentre of the Haar measure on the group of isometries and a Dirac mass $\delta_{\pi}$ at a projection endomorphism $\pi$, then $\xi^{\infty}_{\nu}$ is the average of the isometry-invariant measure and of the Dirac mass on the image of $\pi$. In particular $\xi^{\infty}_{\nu}$ gives positive measure to any hyperplane that contains the image of $\pi$.
@@
Theorem 1.10 implies that we have $E^{+}(\overline{\gamma}_{n})\to l^{\infty}$ almost surely and that $\liminf_{n}\log(\rho_{1}(\overline{\gamma}_{n})/\rho_{2}(\overline{\gamma}_{n})\geq\lambda_{1-2}(\nu)$ almost surely. The question whether this $\liminf$ is actually a limit or not remains open when $\nu$ is not in $\mathrm{L}^{1}$ (See corollary 1.15 for the $\mathrm{L}^{1}$ case). In dimension $2$, we have $\log\frac{\rho_{1}}{\rho_{2}}\leq\log\frac{s_{1}}{s_{2}}$ so we know that $\frac{1}{n}\log\frac{\rho_{1}}{\rho_{2}}(\overline{\gamma}_{n})\to\lambda_{1-2}(\nu)$ almost surely by Theorem 1.7.
@@ 1.5. Regularity results with optimal moment assumptions
The term $Ce^{-\beta n}$ in (11) naturally appears in the proof. This term is annoying because it prevents us from giving a bound on the moments of $\log\frac{|f\overline{\gamma}_{n}v|}{\|f\|\|\overline{\gamma}_{n}\|\|v\|}$. However on can not get rid of it, to give a clear example, let $\nu$ be a lazy measure that gives measure $1/2$ to the identity matrix, then given $fv=0$, we have $\mathbb{P}(f\overline{\gamma}_{n}v=0)\geq 1-2^{-n}$ for all $n$.
@@
Note that $\sum_{k=1}^{+\infty}C\exp(-\beta k)<+\infty$ so $\sum_{k=1}^{+\infty}C\exp(-\beta k)N_{*}{\nu}\left(t/k,+\infty\right)$ goes to 0 as $t\to+\infty$ by the monotonous convergence Theorem. Moreover, since everything is non-negative, we have:
$\mathbb{E}\left(\zeta^{C,\beta}_{\nu}\right)$ $=\int_{0}^{\infty}\zeta^{C,\beta}_{\nu}(t,+\infty)dt$
@@
Given $V\in\mathrm{Gr}(E)$, and $l\in\mathrm{P}(E)$, we write $\mathrm{d}(\mathrm{P}(V),l):=\min_{v\,\in\,V\setminus\{0\}}\mathrm{d}([v],l)$ and given $0<r\leq 1$, we define $\mathcal{N}_{r}(V):=\left\{l\in\mathrm{P}(E)\,\middle|\,\mathrm{d}(\mathrm{P}(V),l)<r\right\}$. Given $p>0$, we say that a random variable $x\in\mathbb{R}_{\geq 0}$ has a finite $\mathrm{L}^{p}$ moment if $\mathbb{E}(x^{p})<+\infty$ and we say that $x$ has a weak $\mathrm{L}^{p}$ moment if $\sup_{t\geq 0}t^{p}\mathbb{P}(x>t)<+\infty$. By integration by parts, we have $\mathbb{E}(x^{p})=\int_{0}^{\infty}pt^{p-1}\mathbb{P}(x>t)dt$ for all $p>0$. So with the same computations as for (14), for all $p>0$, the (weak) $\mathrm{L}^{p}$ moment of $\zeta_{\nu}^{C,\beta}$ is bounded above by the (weak) $\mathrm{L}^{p}$ moment of $N_{*}\nu$ times a constant that depends on $C,\beta$ and $p$.
@@
(16) $\forall V\in\mathrm{Gr}(E)\setminus\{E\},\;\int_{l\in\mathrm{P}(E)}|\log\mathrm{d}(\mathrm{P}(V),l)|^{p}d\xi_{\nu}^{\infty}(l)\leq C^{\prime}.$
@@
Let $l^{\infty}$ be as in Theorem 1.8. We have $l^{\infty}\sim\xi_{\nu}^{\infty}$ so $\xi_{\nu}^{\infty}(\mathcal{N}_{r}(V))=\mathbb{P}(l^{\infty}\in\mathcal{N}_{r}(V))$. Moreover, by Theorem 1.8, we have $\overline{\gamma}_{n}u]\to l^{\infty}$ almost surely and $\mathcal{N}_{r}(V)$ is open, so assuming that $l^{\infty}\in\mathcal{N}_{r}(V)$ implies that there exists $n_{0}$ such that $\overline{\gamma}_{n}u]\in\mathcal{N}_{r}(V)$ for all $n\geq n_{0}$. Therefore, we have $\xi_{\nu}^{\infty}(\mathcal{N}_{r}(V))\leq\lim_{n}\mathbb{P}([\overline{\gamma}_{n}u]\in\mathcal{N}_{r}(V))$.
@@ 2.1. Alignment and contraction coefficient
We remind that we defined the notation $\mathbb{A}^{\varepsilon}$ in Definition 1.3, given $g$ and $h$ two matrices such that the product $gh$ is well defined (i.e., $g$ has as many columns as $h$ has rows) and $0<\varepsilon<1$, we write $g\mathbb{A}^{\varepsilon}h$ as a short notation for $\|gh\|\geq\varepsilon\|g\|\|h\|$. Note that $g$ may have a single line and $h$ may have a single column, in which case $gh$ is a $1\times 1$ square matrix i.e., a scalar. We remind the reader that given $x,y$ two vectors, we have the characterization $\|x\wedge y\|=\min_{a\in\mathbb{K}}\|x-ya\|\|y\|$. Therefore, given $h\in\mathrm{Hom}(E,F)$, we have $\|h\wedge h\|=\max_{x,y\in E\setminus\{0\}}\min_{a\in\mathbb{K}}\frac{\|h(x-ya)\|\|h(y)\|}{\|x\|\|y\|}$.
@@
Note also that for $h$ an endomorphism of rank one, and for all $0<\varepsilon\leq 1$, the cone $U^{\varepsilon}(h)$ is the image of $h$ so it has diameter 0 in the projective space.
@@
The second fact is that $^{t}(gh)=^{t}h^{t}g$. It implies that for all non trivial $g,h$, we have $\frac{\|gh\|}{\|g\|\|h\|}=\frac{\|{{}^{t}h}{{}^{t}g}\|}{\|{{}^{t}h}\|{{}^{t}g}\|}$. The third fact is that $^{t}h\wedge^{t}h=^{t}(h\wedge h)$. It implies that $\sigma(^{t}h)=\sigma(h)$. ∎